OpenStudy (anonymous):

Estimate by rounding to the nearest whole number. 24 2/11 divided by 3 6/7 4 6 8 12 I think its 6 but I want to make sure, I've never been good at fractions.

OpenStudy (anonymous):

\[\frac{ 24 * \frac{ 2 }{ 11 } }{ 3 * \frac{ 6 }{ 7 } }\] Is this the equation? It is difficult to understand your equation by the way you have written it.

OpenStudy (anonymous):

\(\Large 24\frac{2}{11}=\frac{24\times11+2}{11}=\frac{266}{11}\) \(\Large 3\frac{6}{7}=\frac{3\times7+6}{7}=\frac{27}{7}\) \(\Large \Rightarrow \frac{266}{11} \div \frac{27}{7}=\frac{226}{11}\times \frac{7}{27}=\frac{1582}{297}\approx 6\)

OpenStudy (anonymous):

Thank you, and thanks for showing me how you did it to. Fractions confuse me.